OpenStudy (anonymous):

$6,000 dollars is invested in two different accounts earning 3% and 5% interest. At the end of one year, the two accounts earned $220 in interest. How much money was invested at 3%? $2,000 $3,000 $4,000

OpenStudy (anonymous):

i think 2,000

OpenStudy (anonymous):

$4000.

OpenStudy (anonymous):

oh well sorry :p

OpenStudy (anonymous):

\[$4000\times \frac{ 3 }{ 100 }=$120\] \[$2000\times \frac{ 5 }{ 100 }=$100\]

OpenStudy (anonymous):

Total amount invested= $6000 $4000+$2000=$6000 Total interest earned in one year= $220 $120+$100= $220 Therefore, $4000 was invested at 3%p.a. (per annum)